# MyFramelinecolor=blue, outerlinewidth=2pt, roundcorner=20pt, innertopmargin=nerbottommargin=nerrightmargin=20pt, innerleftmargin=20pt, backgroundcolor=gray!50!white

### v1

@@ I Introduction
In this work, we demonstrate through full-wave numerical simulations and quantum computing performance analysis that a space-time-modulated cryogenic-compatible Josephson metasurface significantly enhances the capabilities of superconducting qubit arrays. We present both theoretical and experimental evidence of enhanced connectivity, coherence, and robustness, offering a clear pathway to scalable and high-performance quantum processors.
@@ II.1 Concept of Space-Time-Coupled Qubits
Figure 1: Conceptual depiction of spatiotemporally coupled qubits demonstrating space-time energy transfer and selective excitation of superconducting qubits at distinct quantum states. The process is mediated by a reflective, space-time-periodic cryogenic-compatible superconducting metasurface, enabling precise control over qubit interactions and state transitions to advance quantum information processing and computation.
@@ II.2 Limitation of Conventional Qubit Arrays
Figure 2: Illustration of the restricted functionality in conventional quantum computing qubit arrays, where coupling is limited to monochromatic interactions between adjacent qubits, preventing direct interactions between non-adjacent qubits.
@@ II.4.2 Coherence Time Improvement
The coherence time $T_{2}$ is a critical parameter in quantum computing, as it determines the duration over which a qubit can maintain its quantum state with minimal decoherence. In traditional (monochromatic) qubit systems, $T_{2}$ is often limited by factors such as environmental noise, crosstalk between qubits, and imperfections in control fields. The introduction of a reflective space-time-periodic metasurface enables polychromatic operation, which introduces a frequency separation $\Delta f$ between interacting qubits. This frequency separation mitigates crosstalk, reduces overlap in spectral components, and effectively enhances the coherence time of the system.
The modified coherence time in the polychromatic case is given by:
$T_{2}^{\prime}\approx T_{2}+\frac{\Delta f}{\gamma},$ (5)
@@ II.4.4 Improved Entanglement Robustness
Qubit Fidelity is a measure of the accuracy and reliability of a qubit or a quantum operation. In quantum computing, fidelity quantifies how closely the actual state or operation aligns with the intended or ideal state or operation. High fidelity is crucial for the performance of quantum algorithms, as errors and noise can quickly accumulate and render computations meaningless.
In monochromatic systems, entanglement fidelity is limited by: - Unwanted interactions due to frequency crowding, - Decoherence from shared noise sources.
Fidelity, in the context of quantum mechanics, is a measure of how close two quantum states are. It can be defined for two quantum states $\rho$ and $\sigma$ as
$F(\rho,\sigma)=\left(\text{Tr}\left(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}\right)\right)^{2}$ (7)
@@
By mediating entanglement across multiple frequencies, the metasurface ensures that entanglement fidelity is robust against noise ($F_{\text{entangled}}^{\prime}>F_{\text{entangled}}$). Furthermore, non-adjacent entanglement is feasible, enabling richer quantum states. The entanglement fidelity $F_{\text{entangled}}^{\prime}$ in the polychromatic case is enhanced as:
$F_{\text{entangled}}^{\prime}=F_{\text{entangled}}\cdot\exp\left(-\frac{t}{T_{2}^{\prime}}\right),$ (8)
where the enhanced $T_{2}^{\prime}$ ensures slower decoherence of entangled states.

### v3

@@ II.1 Concept of Space-Time-Coupled Qubits
Figure 1: Schematic representation of space-time-modulated coupling between superconducting qubits, demonstrating energy transfer and selective excitation (e.g., Qubit 1 at state $\hbar\omega_{\text{s}}$ exciting Qubits 2–4 at distinct quantum states $\hbar 2\omega_{\text{s}}$, $\hbar 3\omega_{\text{s}}$ and $\hbar 4\omega_{\text{s}}$). The process is mediated by a reflective, cryogenic-compatible superconducting metasurface with space-time periodicity, enabling multi-frequency control of qubit interactions and selective state transitions to enhance quantum computing efficiency.
@@ II.2 Limitation of Conventional Qubit Arrays
Figure 2: Illustration of the restricted functionality in conventional quantum computing qubit arrays, where coupling is limited to monochromatic interactions between adjacent qubits, preventing direct non-adjacent interactions and rendering the architecture incompatible with arrays of distinct-frequency qubits.
@@ II.4.2 Coherence Time Improvement
The coherence time $T_{2}$ is a critical parameter in quantum computing, as it determines the duration over which a qubit can maintain its quantum state with minimal decoherence. In traditional (monochromatic) qubit systems, $T_{2}$ is often limited by factors such as environmental noise, crosstalk between qubits, and imperfections in control fields. The introduction of a reflective space-time-periodic metasurface enables polychromatic operation, which introduces a frequency separation $\Delta f$ between interacting qubits. This frequency separation mitigates crosstalk, reduces overlap in spectral components, and effectively enhances the coherence time of the system. The modified coherence time in the polychromatic case is given by
$T_{2}^{\prime}=T_{2}\left(1+2\pi\frac{\Delta f}{\gamma}\right),$ (5)
@@ II.4.4 Improved Entanglement Robustness
Qubit Fidelity is a measure of the accuracy and reliability of a qubit or a quantum operation. In quantum computing, fidelity quantifies how closely the actual state or operation aligns with the intended or ideal state or operation. High fidelity is crucial for the performance of quantum algorithms, as errors and noise can quickly accumulate and render computations meaningless. In monochromatic systems, entanglement fidelity is limited by unwanted interactions due to frequency crowding, and decoherence from shared noise sources. Fidelity, in the context of quantum mechanics, is a measure of how close two quantum states are. It can be defined for two quantum states $\rho$ and $\sigma$ as
$F(\rho,\sigma)=\left(\text{Tr}\left(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}\right)\right)^{2}.$ (7a)
@@
The metasurface’s polychromatic coupling architecture enhances entanglement robustness by mediating multi-frequency interactions. This approach suppresses noise-induced decoherence while enabling direct entanglement between non-adjacent qubits, thereby expanding the Hilbert space accessible for complex entangled states. Under Markovian decoherence [38, 39]—a regime applicable to superconducting qubits dominated by short-correlated noise—the entanglement fidelity $F_{\text{entangled}}^{\prime}$ in the polychromatic case reads
$F_{\text{entangled}}^{\prime}=F_{\text{entangled}}\cdot\exp\left(-\frac{t}{T_{2}^{\prime}}\right),$ (7b)
which is the enhanced coherence time enabled by frequency separation $\Delta f$. The exponential fidelity decay slows proportionally to $T_{2}^{\prime}$, ensuring $F_{\text{entangled}}^{\prime}>F_{\text{entangled}}$ for identical noise conditions. This fidelity improvement arises from the metasurface’s spectral isolation of qubit interactions, which reduces overlap with broadband noise and suppresses correlated errors in multi-qubit operations.